OpenStudy (anonymous):

Find the length of the rectangle whose vertices are given. A(2, 1), B(5, 4), C(0, 3), D(3, 6)

OpenStudy (anonymous):

Length of rectangle? Or Area or perimeter?

OpenStudy (anonymous):

In order to find the lengths of the rectangle, use distance formula and find the distances between AB, BC, CD and AD.

OpenStudy (anonymous):

So... \[l(AB)=\sqrt{(2-5)^2+(1-4)^2}=\sqrt{(-3)^2+(-3)^2}=\sqrt{9+9}=\sqrt{18}=3\sqrt{2}\] \[l(BC)=\sqrt{(5-0)^2+(4-3)^2}=\sqrt{(5)^2+(1)^2}=\sqrt{25+1}=\sqrt{26}\] \[l(CD)=\sqrt{(0-3)^2+(3-6)^2}=\sqrt{(-3)^2+(-3)^2}=\sqrt{9+9}=\sqrt{18}=3\sqrt{2}\] \[l(AD)=\sqrt{(2-3)^2+(1-6)^2}=\sqrt{(-1)^2+(-5)^2}=\sqrt{1+25}=\sqrt{26}\] From Above results we can see that l(AB)=l(ACD) & l(BC)=l(AD) i.e. opposite sides are congruent